OpenStudy (anonymous):

Find the unit rate. forty books in sixteen minutes ten books in 4 minutes forty books in 1 minute two point five books in 1 minute 1 book in sixteen minutes

OpenStudy (texaschic101):

16/40 = x/10 40x = 160 x = 4 answer is 10 books in 4 minutes